this is very difficult for me, can you help?
Let f(x)=5x-6 and g(x)=x^2-8x+16. find the following:
1) (f+g)(x)
2) (g-f)(2)
3) (fg)(5)
4) f/g(x)
thanks so much!

Answer by stanbon(75887)

Let f(x)=5x-6 and g(x)=x^2-8x+16. find the following:
1) (f+g)(x)
=(5x-6)+(x^2-8x+16)
=x^2-3x+10
--------------
2) (g-f)(2)
=[(x^2-8x+16)-(5x-6)](2)
=[x^2-13x+22](2)
=2^2-13(2)+22
=4-26+22
=0
--------------
3) (fg)(5)
=(5x-6)(x^2-8x+16)
=5x^3-40x^2-6x^2+80x+48x-96
=5x^3-46x^2+128x-96
Evaluate at x=5 to get:
=5(5^3)-46(25)+128(5)-96
=19
-------------
4) f/g(x)
=(5x-6)/(x^2-8x+16)
---------
Cheers,
Stan H.